OpenStudy (anonymous):

z-5 to the 2 power for z =6? show work

hero (hero):

1

hero (hero):

\[(z-5)^2 = (6-5)^2 = 1^2 = 1\]

jimthompson5910 (jim_thompson5910):

\[(z-5)^2\] \[(6-5)^2\] \[(1)^2\] \[1\]So \[(z-5)^2=1\] when z=6

OpenStudy (saifoo.khan):

(z-5)^2 (6-5)^2 1^2 1

hero (hero):

My solution takes up less space